A 50 kg object moves at 50 m/s around a circular path of radius 10m. What centripetal force is needed to keep it in its orbit? What is the centripetal acceleration of the object?


1
Expert's answer
2021-12-13T16:56:30-0500

The centripetal force needed to keep it in its orbit is


"F_c=mv^2/r=12500\text{ N}."

The centripetal acceleration:


"a_c=v^2/r=250\text{ m/s}^2."